### v2

# Visual-auditory Extrinsic Contact Estimation

Abstract
Robust manipulation often hinges on a robot’s ability to perceive extrinsic contacts—contacts between a grasped object and its surrounding environment. However, these contacts are difficult to observe through vision alone due to occlusions, limited resolution, and ambiguous near-contact states. In this paper, we propose a visual-auditory method for extrinsic contact estimation that integrates global scene information from vision with local contact cues obtained through active audio sensing. Our approach equips a robotic gripper with contact microphones and conduction speakers, enabling the system to emit and receive acoustic signals through the grasped object to detect external contacts. We train our perception pipeline entirely in simulation and zero-shot transfer to the real-world. To bridge the sim-to-real gap, we introduce a real-to-sim audio hallucination technique, injecting real-world audio samples into simulated scenes with ground-truth contact labels. The resulting multimodal model accurately estimates both the location and size of extrinsic contacts across a range of cluttered and occluded scenarios. We further demonstrate that explicit contact prediction significantly improves policy learning for downstream contact-rich manipulation tasks.
Keywords: Multimodal Perception, Contact-Rich Manipulation, Sim2Real
Figure 1: (a) Our proposed fingers with an active conduction speaker and contact microphone emitting and receiving sound through the object. The absorption and reflection of the audio from the contact between the object and environment enables extrinsic contact estimation despite visual ambiguities. Challenges include (b) where objects occlude the contact between the box and the table, (c) where a different surface type can change the acoustic feedback, and (d) estimating the object’s contact status for near-contact scenarios.
footnotetext: * Denotes equal contribution.footnotetext: Correspondence to: {yixili, jayjun}@umich.edu.
1 Introduction
Extrinsic contact estimation is a crucial capability for robots to accurately understand how tools interact with their environment. Properly perceiving these contacts enables the robot to plan and control its actions effectively. Vision-based methods allow observation of the entire scene, but they often fall short in providing sufficient local information, particularly when contacts are occluded or within the resolution limits of the sensor. While tactile sensors or force/torque sensors offer precise measurements of direct contact surfaces, they struggle to perceive indirect contacts, such as those between a tool and the environment. This challenge creates a sensing gap in extrinsic contact estimation. As illustrated in Fig. 1(d), distinguishing whether the objects are in contact when they are close to each other within the sensor’s resolution is difficult. Similarly, in Fig. 1(b), occlusions make it challenging to determine contact status as it is not directly observable.
To address this challenge, we propose a novel visual-auditory extrinsic contact estimation method, VA2Contact. Our approach integrates global visual feedback with local information obtained through active audio sensing, as illustrated in Fig. 1(a). This enables the estimation of extrinsic contacts as masks in 2D image space, from which both the contact location and type can be inferred. Additionally, we introduce an innovative audio-hallucination technique to overcome the difficulty of scalably obtaining audio feedback in simulation. This technique involves injecting real-world audio data into the simulation dataset with corresponding labels, thereby bridging the sim2real gap.
Our work builds upon and improves recent advancements in the field. A closely related study is Im2Contact [1], where a single RGB-D camera was used to learn a probability map of contact from depth images, optical flow extracted from RGB images, and proprioception. While they demonstrated a promising ability to estimate contact locations in pixel space, two main issues persist. The first is that the existing method struggles with obtaining the location and shape of the contact patch in static scenes as it relies solely on the optical flow for extracting the temporal information of movement in the scene to infer contacts and such indication of motion is not always correlated to contact events. Second, the existing approach is limited by occlusions and the resolution of the camera. By incorporating local active-audio information, our approach provides more comprehensive contact details, regardless of the contact type, and in spite of heavy occlusion, addressing these limitations and enhancing the performance of binary extrinsic contact detection and the accuracy and robustness of estimating the geometry of the extrinsic contact patches.
2 Related Works
Extrinsic Dexterity and Contact Estimation: In the field of manipulation, external contact sensing plays a crucial role. “Intrinsic contact,” which refers to the direct contact between the robot and the environment, has been well studied [2, 3, 4]. However, when manipulating tools, estimating “extrinsic contact,” such as sensing the contact between a tool held by the robot and the environment, becomes important. Sensing extrinsic contact is more challenging due to the indirect transmission of contact force/torque and the uncertainties in the object’s geometry, stiffness, and pose. [5] combines neural fields and vision-based tactile sensing to estimate the probability of extrinsic contacts between object and environment for any point on the object’s surface. Most studies rely on force/torque sensors or tactile sensors, which often involve strong assumptions or are limited to predefined contact configurations [6, 7, 8]. Some prior works require knowledge about the tool and the environment. For instance, [9] and the subsequent work [10] focused on estimating object pose and contact simultaneously from force/torque feedback. Additionally, data-driven methods have shown potential in estimating extrinsic contacts. [1] demonstrated that with visual data and robot proprioception, their model could predict the contact location in image space without assuming prior knowledge about the object and environment. [11] uses active tactile feedback to regulate a more consistent contact mode and make contact estimates for peg insertion tasks. Recently, multimodal policies that combine proprioception, vision, and audio have also been explored to decompose tasks into stages [12] and by finetuning vision-language-action models [13] for contact-rich manipulation tasks.
Audio for Robotic Manipulation: Audio is also a widely used modality in the field of robotics. Due to its physical principles, audio signals can provide frequency-related characteristics. Passive acoustic sensing directly uses sound waves from structural vibrations [14]. Some studies have modeled the sounds of object-surface interactions [15]. This approach is also commonly used in soft pneumatic actuators (SPAs) to sense state changes. Previous studies also showed that in end-to-end robot learning algorithms, including audio signals as input improved task performance [16, 17, 18, 19]. SonicSense [20] recently introduced a method to extract object material and shape information from in-hand acoustic vibrations, highlighting how such passive audio cues can enhance object understanding and contact estimation during manipulation. Other works also use active audio sensing methods to gain more information from the environment. Zöller et al. detected contact by embedding a microphone into an SPA to measure the sound induced by contact [21]. Similarly, studies [22, 23, 24, 25] proposed embedding a microphone and speaker in an SPA, playing sweeping sounds, and measuring changes from the microphone to sense deformation or contacts. Multimodal sensing with audio can be used for surface proximity detection with piezoelectric transducers for robot collision avoidance [26]. There are also studies using active audio on rigid manipulator grippers to estimate grasping positions and for object recognition [27]. [28] uses a microphone attached at the gripper to receive audio feedback on contact events under partial observability.
In this work, we leverage acoustic signals that transmit through solid objects by combining this with vision-based methods to estimate extrinsic contacts using a model-free approach. Unlike prior approaches that rely on fixed sensor setups or require assumptions about contact geometry, our method generalizes across contact types and tool configurations by fusing sound and vision in a flexible learning pipeline. Moreover, we introduce active-audio sensing to probe extrinsic contacts.
3 Methodology
Figure 2: System Architecture. Visual-Auditory Extrinsic Contact Estimation (VA2Contact) is trained in simulation with real-world audio collected through our active-audio sensing mechanism. Raw audio waveforms are processed with Short-Time Fourier Transform (STFT). VA2Contact can zero-shot transfer to the real-world for contact prediction tasks. Here, the output contact probability map is overlaid onto the full depth image of the scene. Note the usage of off-the-shelf metric depth estimation models (Depth-Pro [29]) and optical flow estimation model (RAFT [30]) both for scalable sim-based training and real-world inference, to bridge the sim-to-real gap effectively. VA2Contact unlocks contact perception under occlusions for contact-rich tool manipulation tasks such as wiping, which we demonstrate through real-world policy learning experiments.
Problem Statement: Consider a robot holding an unknown object and making contact with an unstructured environment. The goal of our method is to estimate the the location and shape of the extrinsic contact between the grasped object and the environment. We assume no prior knowledge about the grasped object or environment. The inputs to our method are a depth map from a statically mounted camera, an optical flow image, fixed-length audio signatures measured at gripper finger tip, and proprioceptive sensing of the robot state.
Method Overview: We train a multimodal model that can predict a per-pixel contact probability map over the scene given the inputs. All the training data except for audio are synthetically generated in simulation where pre-recorded audio data from the real-world are injected. The model is then zero-shot transferred to the real-world as in Figure 2. In the following, we outline the active-audio sensing mechanism that enables extrinsic contact sensing beyond vision, data generation process to train such a model, the model architecture, and training details unique to this problem.
Figure 3: Active-audio Sensing. (1). A sweeping acoustic signal is generated from (2) using conduction speaker finger, where (3) the sound propagates through the object and vibrates with any extrinsic contact it makes, and (4) which is received at the contact microphone finger, and (5) the audio waveform is converted to a spectrogram.
Active-Audio Sensing: Suppose the robot is dragging the object it is grasping across a surface. The physical interaction between the object and environment creates an audio signature that contains a wealth of information about the interaction including its extrinsic contact patch size/shape, object/surface material types, contact location, and movement speed. However, if the object is in a static contact with the environment, there will be no audio feedback from the relative motion of the object, as audio signals require a source, either motion or an active audio source. To overcome this limitation and generalize to such static scenes, we introduce an active-audio system where one finger acts as an actuator (i.e. a speaker) by emitting controlled acoustic signals (i.e. a sweeping impulse signal), while the other finger as a receptor (i.e. a microphone) that receives the acoustic feedback, which changes depending on the grasped object’s material properties and the presence of extrinsic contacts. As illustrated in Figure 3, contact with the environment alters the received signal due to energy absorption, enabling active contact perception even in the absence of motion. This mechanism extends the robot’s perception to static contact scenarios, complementing vision and proprioception. Even without motion, contact with the environment alters the acoustic response by absorbing energy and modifying the transmission characteristics. This allows the system to distinguish between contact and non-contact scenarios based on the presence and properties of the received audio signal.
Challenges in Data Generation: However, one major challenge in training a multimodal model for the task of visual-auditory extrinsic contact estimation is obtaining a dataset with aligned vision, audio, and ground-truth contact patch labels. In the real-world, it is relatively easy to obtain the audio information along side with both depth maps and robot proprioception data. However, obtaining correct contact patch masks is not only hard but also inaccurate with human labels, because there is no direct way to observe the exact contact patch shape and size. Meanwhile in simulation, it is simple to generate a large dataset with both correct contact patch masks and correct labels. However, obtaining realistic audio signals from simulation remains a great challenge.
Real-to-Sim Audio Hallucination: To overcome this challenge, we propose an audio hallucination technique. In the real-world, we collect audio signals ($\mathbf{v}_{t}$) as the robot manipulating various grasped objects by making diverse contact with the environment. During the robot motion, we record contact and contact geometry labels ($\mathbf{g}_{t}$) from the set [free, point, line, patch]. We emphasize that we do not record contact locations, only the simple-to-observe contact labels. Thus the labeled dataset contains audio samples paired with contact types, represented as $\mathcal{D}_{\text{audio}}=\{(\mathbf{v}_{0},\mathbf{g}_{0}),\cdots,(\mathbf{v}_{n},\mathbf{g}_{n})\}$. Next, in simulation we create scenes with the same robot and table as well as a large distribution of objects that the robot can grasp and interact with. This distribution of objects is not the same as the test objects used in the real-world setup. From the simulated visual scene, we obtain a depth map ($\mathbf{d}_{t}$) and obtain proprioceptive state ($\mathbf{p}_{t}$). We also take one depth map when the object is lifted without contact as a reference frame ($\mathbf{r}_{t}$) of the grasped object. Although the simulator can generate optical flow images, we generate the optical flow ($\mathbf{f}_{t}$) using an off-the-shelf RAFT model [30] to later facilitate sim-to-real transfer, where we use the same model to generate the optical flow images for the real world test data and similarly use Depth-Pro [29] for depth. It is noted that we generate optical flows only from depth maps to avoid the influence from shadows that would break the color constancy assumption. Most importantly, simulation also directly provides us with the contact geometry label and its corresponding shape ($\mathbf{s}_{t}$) when contact occurs. At each contact event, we randomly select a sample of the real-world audio corresponding to the label provided by the simulator and create a labeled dataset of contacts $\mathcal{D}=\{(\mathbf{O}_{0},\mathbf{s}_{0}),\cdots,(\mathbf{O}_{m},\mathbf{s}_{m})\}$ where the first element is the observation vector $\mathbf{O}_{i}=(\mathbf{v}_{i},\mathbf{d}_{i},\mathbf{r}_{i},\mathbf{f}_{i},\mathbf{p}_{i})$ and the second is the label.
The premise of our approach is that, while audio signatures vary across specific object interactions, they share consistent patterns that reflect the underlying contact geometry. A model trained on such data can learn to generalize by ignoring object-specific details and focusing on features indicative of contact type. For instance, point or line contacts tend to produce higher-pressure interactions, resulting in sharper, high-frequency sounds. In contrast, patch contacts generate a more diffuse sounds resembling white noise, with energy spread across a wider frequency range due to a larger contact area. A brief illustrative case study is provided in Appendix B using power spectral analysis.
Model Architecture: For visual-auditory extrinsic contact estimation, our model architecture builds on the UNet architecture [31] as illustrated in Figure 2(b). The model consists of three streams of UNet encoders, each processing different streams of data. The first stream encodes a cropped depth image of the object at a reference frame. The second stream encodes a depth image and an optical flow image at the current frame, which are cropped around the projected end-effector pixel coordinates and stacked along the channel dimension. The third stream encodes an image of a log-mel spectrogram from a 1s long audio waveform sampled at 44.1 kHz. All input images are of shape $256\times 256$ and the 7-DoF robot pose is fused at the bottleneck of the UNet.
Additionally, a prior work [1] suggests a few strategies in which we adopt to facilitate better generalization and sim2real transfer. First, cropping the input images of the task scene centered around the end-effector position projected onto the pixel space simplifies the learning problem by removing background distractions and with an improved focus on the tool-environment interaction. For every cropping operation, additional three channels are stacked for coordinate convolution [32] to provide the model with an explicit information about the spatial location of each pixel within the cropped image and help the model maintain the understanding of spatial relationships even after cropping. Another major challenge with the extrinsic contact estimation is when the grasped tool is heavily occluded in a cluttered scene that the parts of the grasped object that make contact with the environment is not visually observable. Thus, we opt to provide the model with a contact-free reference depth image of the object in the gripper. Moreover, visual ambiguities are present when differentiating an in-contact versus a near-contact state of the robot. This can be partially resolved by introducing the optical flow as an input to represent the temporal information of the robot motion.
We further utilize the audio modality to acquire active feedback from extrinsic contact and interaction with the environment from the sine sweep impulse response that changes per contact mode type (refer to Appendix B), grasped object, and robot motion. The output of the model is a contact probability map that can be thresholded to obtain the contact patch as a single channel mask image.
4 Implementation
Active-Audio Processing: The frequency characteristics of each object and contact vary depending on their physical properties, including stiffness and geometry. To capture these variations comprehensively, we transmit a sweeping impulse sound that spans a broad frequency range (20 to 20k Hz) through the object repeatedly at 1 Hz. We collect the audio feedback from a microphone and then use Short-Time Fourier Transform (STFT) and scaling to generate a log-mel spectrogram using 64 mel frequency bins. In the spectrogram, shown in Figure 3, the x-axis represents time, the y-axis represents frequency, and the color represents energy at that time and frequency. The frequency axis is non-linearly scaled to show more information in the lower frequency range, which typically contains more useful information for contact. Given its ability to display both temporal and frequency information and their relationship, we choose this as the audio representation. We fix the duration of each audio sample to one second, ensuring that it captures the full impulse response across the entire sweeping frequency range.
Real World Setup: For collecting and generating such scalable dataset for extrinsic contact perception, we use the following setup. In both simulation and the real-world, we use a 7-DoF Franka Emika Panda robot arm in a tabletop environment. Active-audio fingertips, illustrated in Fig. 3 are mounted onto the gripper, and both input and output audio signals are routed to a computer through a Focusrite Scarlett 4i4 3rd Gen amplifier. We set a single RealSense RGB-D camera pointing to the table for collecting visual feedback. We selected several objects from YCB dataset [33] (a spoon, clamp, tube, apple, lemon, orange, pear, screw driver, scissors, pods) and a few other objects such as a RealSense box and a dustpan as test objects. Note that these are not seen during training.
Real World Audio Collection: To collect the active audio feedback with different types of contact, after the robot grasps an object, we teleoperate the robot to random orientation, and make contact with the environment. During the execution of each motion, we maintain the same contact mode(free, point, line, or patch contact), we also collect the active audio feedback from the scene, with sample rate of 44100 Hz. To provide the active audio signal, we input sweeping sound from 20Hz to 20000Hz in 1 second, and repeat it at 1 Hz. We also label the audio given the contact mode. We collected 1000 audio samples with variety of objects, contact shape and contact mode.
Simulation Setup and Real-to-Sim Audio Injection: We use the GPU-based Isaac Gym [34] simulation to replicate the real-world scene generating diverse extrinsic contact data at scale. In each episode, the robot grasps random objects and rotate to a random orientation, then lower down to the tabletop, and starts sliding the object on the table. For each environment, we randomly select cubes with random size and objects from ycb dataset[33] that are smaller than 10cm in diameter. We obtain the contact points of the object and the environment, then project them back to the camera frame to obtain the contact masks. We also obtained a contact label from the size and shape of the contact, which are [free, point, line, patch]. With these labels, we randomly select real-world audio samples with the same labels and then pair with depth image. The training data contains 6880 episodes where each episode contains around 60 frames, which approximates to a total of 440k data samples.
Figure 4: Different contact modes learned from simulation data. The cropped depth is centered around the projected pixel coordinate of the EE pose. Sample predictions from test simulation data are shown as contact probability maps overlaid on top of scene depth.
Training Details: We train our model using the pixel-wise binary cross entropy loss with logits using the ground-truth contact masks obtained from simulation. We train for 20 epochs with a learning rate of $45e-$ 4 on a single A6000 GPU.
Real World Test Data Collection: To evaluate VA2Contact on extrinsic contact detection, we collect data that consists of RGB-D images of the scene, EE pose, camera extrinsics, a reference RGB-D image, audio feedback, and a manually labelled extrinsic contact mask. We use the same off-the-shelf models (Depth-Pro and RAFT) to obtain consistent depth and optical flow images
Additionally, for collecting the real-world depth images, we primarily use the monocular metric depth estimation model Depth-Pro [29] to generate hole-less depth map from raw RGB images of the scene (refer to Appendix A to see noisy depth images that are significantly out-of-distribution). We collected data in 4 cases: general, occlusion, near-contact, and different surface cases. We use RAFT [30] to generate optical flow images. We use the same audio collection pipleline as audio collection for test data collection.
For downstream policy learning with VA2Contact in the perception pipeline, we use an Oculus Quest 2 device to teleoperate and collect 41 demonstrations for the wiping task that amount to 16k timesteps of training data with audio feedback at 15 Hz. We preprocess this data to generate consistent depth, optical flow, and contact maps using Depth-Pro, RAFT, and VA2Contact to train diffusion policy that takes two images (one for raw RGB and one for contact prediction image) and joint state as input. The diffusion policy is trained on relative joint action space that predicts a action horizon of 16 where we execute the first 8 for receding horizon control at 15 Hz.
5 Experiments
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 0.70 0.94 0.81 0.69 0.023 0.058 19.2
VA2Contact w/o flow 0.98 0.95 0.96 0.95 0.030 0.073 13.6
VA2Contact 0.86 0.90 0.88 0.83 0.027 0.28 4.8
Table 1: Real world results for sim-to-real extrinsic contact estimation on S1 to test on general contact cases. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
5.1 Extrinsic Contact Estimation
Baselines and Ablations: We compare our visual-auditory method against a similar model-free and vision-only method Im2Contact [1]. We also evaluate a variant of VA2Contact trained without optical flow to experiment if audio can fully replace it.
Real World Test Scenarios: To evaluate the sim2real transfer of our models on the extrinsic contact estimation problem, we design the following real world test scenarios. These scenarios highlight the primary challenges, including demonstrating sim2real transfer under substantial sim2real gap (S1, S2), contacts invisible to vision (S3, S4).
Figure 5: Sim-to-Real Transfer and Real-world Extrinsic Contact Predictions for S1. The RGB-D images with GT contact probability masks are wrapped in red. Three models, VA2Contact, VA2Contact w/o optical flow, and Im2Contact, are tested where the results are shown per column. VA2Contact’s contact probability predictions are overlaid to depth images, wrapped in green. The color bar represents contact probability (0.0 $\leftarrow$ | $\rightarrow$ 1.0). All grasped objects used for real-world testing are unseen geometries (a cup, pear, dustpan, box, blue sponge, lemon, can, clamp). VA2Contact is able to zero-shot predict diverse contact types over objects with varying properties.
  • S1: Generalization to unseen gripper-held object types and geometry. (Table 1 and Figure 9)
  • S2: Generalization to unseen table surface to introduce audio variation. (Appendix C)
  • S3: Near contact cases that are not in contact. (Appendix C)
  • S4: Cases where the contact between gripper-held object and the table is occluded. (Appendix C)
Experimental Results: The main results of the sim2real transfer of visual-auditory extrinsic contact estimation in the real world setup are shown in Table 1 and Figure 9. We further explore applications to edge cases such as visual occlusions of the contact and near-contact setups where the active-acoustic signal could provide important cues. In the general case, the model with audio overall exhibits higher recall and F1 score, indicating a superior ability to detect true contacts. This suggests that audio cues help reduce false negatives without substantially increasing false positives. VA2Contact w/o optical flow scores the highest across all metrics for binary contact detection because audio provides a more direct information about contact whereas optical flow just tells us where things moved, which might not be directly correlated. However, in terms of the actual contact patch prediction, VA2Contact which uses both active-audio and optical flow performs the best as shown in Figure 9. For other scenarios (S2-4) where an additional sensor modality can be useful, we observe that VA2Contact does improve contact perception under heavy visual ambiguities. For more details and predictions, refer to Appendix C.
5.2 Contact-aware Policy Learning
Figure 6: Real-world wiping task scene with contact-aware diffusion policy, enabled by VA2Contact.
Given this sim2real visual-auditory extrinsic contact sensing capability, we evaluate whether this can improve the downstream policy learning and performance for contact-rich and occlusion-heavy manipulation tasks. We selected a wiping task, which requires the robot to hold a marker eraser and maintain contact to erase the markers. To test this, we train a baseline diffusion policy with just an RGB image stream and a contact-aware diffusion policy as in Figure 2 that takes the contact probability map image in addition to the RGB image stream as inputs. We evaluate the performances over 10 rollouts for each model. The baseline model achieved 4/10 success due to inconsistent contact with the whiteboard whereas the contact-aware policy achieved 8/10 success rate.
Overall, the integration of audio input into the contact estimation model enhances its ability to reliably detect and localize contacts in complex real-world environments and bridges the sim-to-real transfer well. The audio cues effectively complement visual data, particularly in situations where visual information is incomplete or ambiguous such as under occlusions, which are common in real world manipulation tasks. Moreover, we find that such multimodal contact-aware representations can enhance contact-rich manipulation capabilites further.
6 Conclusion and Discussions
Summary: We presented a method trained entirely in simulation that accurately estimates extrinsic contacts by leveraging multimodal signals—fingertip audio, vision, and proprioception—without relying on prior knowledge of the object or environment. Active audio sensing is shown to provide rich, local information about contact events and modes, effectively complementing vision in contact-rich manipulation tasks. By injecting realistic audio signals into simulation data, we bypass the difficulty of simulating audio while still improving contact prediction performance. Our method demonstrates strong sim-to-real transfer, especially in visually ambiguous contact scenarios, outperforming baselines that exclude audio input.
Limitations: Despite its effectiveness, our system faces several limitations. Internal vibrations from the robot—especially within the Franka gripper—introduce significant noise, reducing the signal-to-noise ratio (SNR) and impacting audio quality. Improving vibration isolation at the fingertip could substantially enhance signal clarity and sensing robustness.
The contact masks in the real world test dataset are annotated manually, with the help of RGB image collected along with depth. It is noted that these masks are subject to the annotator’s subjective experience due to the occlusion and difficulty to directly observe the ground truth contact.
The current audio representation, while effective, may not capture the full range of contact features. Exploring alternative encodings or integrating large pretrained audio models could yield more expressive and robust features for contact reasoning.
Our dataset focuses on rigid objects with low vibration absorption to ensure clean audio, but this restricts generalizability. Softer, more deformable objects and liquid pose challenges due to annotation complexity and high signal variability, yet incorporating them is essential for broader applicability.
In addition, we have yet to fully explore contact chains, where signals travel across multiple interfaces and suffer cumulative loss. Addressing these scenarios will require more sensitive sensing, improved actuation—such as the potential use of vibration motors—and careful attention to SNR across all system components.
Appendix A Real World Depth Images
The figure below displays how noisy depth images can be despite using a higher-end high resolution 3D cameras compared to a simulation environment that returns an optimal depth maps.
Figure 7: A comparison between a depth image from simulation and a raw and a sample processed depth images from Photoneo 3D scanner. The real depth images present a significant distribution shift from the simulation training data.
Appendix B Power Spectral Analysis of Acoustic Signatures across Contact Modes
As illustrated in Figure 8, each contact mode exhibits distinct characteristics in its power spectral density (PSD), revealing informative patterns tied to contact geometry. In the low-frequency range (0–100 Hz), all modes show a steep initial drop in spectral power, with the patch contact displaying the lowest overall magnitude—indicative of more dampened or lower-energy vibrations. The first peak, observed around 100–300 Hz, is both wider and higher for the point contact compared to line and patch contacts. This reflects the sharper, higher-energy acoustic responses produced by the small contact area and high contact pressure of point interactions. In the mid-frequency range (300–1000 Hz), the line contact maintains higher PSD values than the point contact, which itself remains consistently above the patch. This suggests that line contacts produce more sustained vibrational energy across a broader range of frequencies. In the high-frequency region beyond 1000 Hz, all contact types show a gradual convergence, though the line contact still dominates in PSD magnitude, and the patch contact continues to exhibit the lowest profile. Overall, these trends indicate that the size and type of contact area—point, line, or patch—have a significant impact on the spectral shape of the resulting acoustic signal. The presence of consistent and structured spectral differences highlights the potential for learning-based models to classify contact modes based on audio features alone.
As our method utilizes the contact mode type labels to inject audio into the simulated episodes of contact, we analyze the frequency components of the different contact modes using various household objects and different surface types to showcase how each mode displays similar features despite variances in vibrational properties, material type, rigidity, and content. We recorded 3s long audio waveforms where a sliding contact is made between each object and surface of the environment. However, it is noted that it is difficult to exert a consistent force over different setups so we collect three audio data per setup and average. We collected the audio signals over four object types (a box, an empty plastic bottle, a power strip, a metallic screw driver set case) on two different surface types (a powder coated steel and a wooden surface) and analyzed the power spectral density averaged across different contact modes. The acoustic signal features and patterns can be explained through the contact pressure and the contact size it makes with the environment. Assuming the same force is being exerted, the contact pressure is highest from the point, line, and patch, in this order as the contact area increases. The patch contact in general has a relatively horizontal profile after the steep drop in the low frequency range from 0 to 100 Hz. In contrast, shifting focus to the first peak of the point and line contact plots, we see slight difference in the bandwidth of the peak where the point contact is wider and higher. As the frequency increases the point contact plot maintains a profile that stays under the line contact’s profile. It should be noted that the forced applied at different contact modes is not consistent and there are more variety in combinations possible in the real world scenarios. However, this figure presents that there exists a learnable pattern in which audio signal may be from which contact mode.
Figure 8: Averaged power spectral density (PSD) of acoustic feedback from sliding contact interactions across three contact modes: point, line, and patch. Point contact exhibits a broad, high-amplitude peak in the 100–300 Hz range, line contact shows elevated energy across mid to high frequencies, and patch contact presents a flatter, more damped spectral profile. Results are averaged over four objects and two surface types, highlighting the consistent, mode-specific spectral signatures in contact-generated audio.
Appendix C More Results for Extrinsic Contact Estimation
Different Surface Type Cases (S2): For scenarios involving a different surface in Table 2, both models scores high in binary contact detection metrics. For the contact geometry estimation, the performance of the two models are similar for the model with audio or even better for the model without audio in terms of IOU. This indicates that while our method can transfer to different surface types, which might lead to a variation in audio feedback. The model without audio tends to be less affected by this change as it solely reasons through vision. This indicates that while visual cues are sufficient for high precision, audio input can be generalized across varying surfaces with consistent contact detection.
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 1.00 0.78 0.88 0.78 0.04 0.139 10.52
VA2Contact w/o flow 1.00 0.86 0.92 0.86 0.060 0.116 10.74
VA2Contact 1.00 0.94 0.97 0.94 0.054 0.221 4.36
Table 2: Real world results for sim-to-real extrinsic contact estimation on S2 for Evaluating on Different Surface Types. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
Near-contact Cases (S3): In the near contact scenario in Table 3, where no ground truth contact masks are available as there is no contact event, models that have audio significantly outperform the model that relies solely on vision. Most of Im2Contact’s predictions are false positives, which indicates the inability to disambiguate near contact cases. Yet our VA2Contact can effectively probe with active-audio sensing to differentiate being in contact and not in contact despite visual ambiguity.
Models Binary Contact Detection
TP $\uparrow$ TN $\uparrow$ FP $\downarrow$ FN $\downarrow$
Im2Contact 0 2 48 0
VA2Contact w/o flow 0 48 2 0
VA2Contact 0 49 1 0
Table 3: Real world results for sim-to-real extrinsic contact estimation on S3 for Evaluating on Near-Contact Cases where there is no actual contact being made. Models are evaluated on binary contact detection with TP (true positive), TN (true negative), FP (false positive), FN (false negative).
Occlusion Cases (S4): Under occlusion in Table 4, the benefits of audio are even more pronounced that the VA2Contact with audio outperforms Im2Contact. Our model with audio achieves a recall of 0.94 compared to 0.78 without audio, which indicates VA2Contact’s ability to better perceive contacts under occlusions, and an F1 score of 0.97 versus 0.88. The accuracy also improves significantly (0.94 vs. 0.78). These improvements highlight the model’s enhanced capability to detect contacts that are visually obstructed, leveraging audio cues to compensate for the lack of visual information.
Models Binary Contact Detection Contact Location, Shape, & Size
Prsn $\uparrow$ Rcll $\uparrow$ F1 $\uparrow$ Acc $\uparrow$ BCE $\downarrow$ IOU $\uparrow$ CD ($\text{px}^{2}$) $\downarrow$
Im2Contact 0.91 0.89 0.90 0.82 0.029 0.079 15.3
VA2Contact w/o flow 1.00 0.87 0.93 0.88 0.025 0.099 10.84
VA2Contact 0.96 1.00 0.98 0.96 0.034 0.22 10.16
Table 4: Real world results for sim-to-real extrinsic contact estimation on S4 for Occlusions. Models are evaluated on binary contact detection and the accuracy of contact geometry estimation, which is only provided for the true positive cases where the model predicts a mask above a threshold level and the ground-truth is in-contact. CD: Chamfer Distance. IOU: Intersection over Union. BCE: Binary cross-entropy.
Figure 9: Sim-to-Real Transfer and Real-world Extrinsic Contact Predictions for S2 and S4. The RGB-D images with GT contact probability masks are wrapped in red. Three models, VA2Contact, VA2Contact w/o optical flow, and Im2Contact, are tested where the results are shown per column. VA2Contact’s contact probability predictions are overlaid to depth images, wrapped in green. The color bar represents contact probability (0.0 $\leftarrow$ | $\rightarrow$ 1.0). All grasped objects used for real-world testing are unseen geometries (a cup, pear, dustpan, box, blue sponge, lemon, can, clamp). VA2Contact is able to zero-shot predict diverse contact types over objects with varying properties.